6. Consider a voltaic cell based on the following cell reaction:
Ni(s) + At2(s) ? Ni2+(aq) + 2At– (aq)
Given that the standard cell emf is 0.55 V, what is the standard reduction potential for astatine? [E°(Ni2+/Ni) = –0.25 V]
Answer #1

E0(cell) = E0(At2/At-) - E0(Ni2+/Ni)

0.55 = E0(At2/At-) - (-0.25)

E0(At2/At-) = +0.30 V